In an oil exploration in Arabic sea, suppose that the probability of an oil strike is 1 in 500 drilling. What is the probability of having exactly 1 oil producing well in 800 explorations.



1
Expert's answer
2022-04-15T14:56:58-0400

Let p = 1 / 500 = 0.002, n = 800

According to the binomial distribution:

"P(x=k)=C_n^kp^k(1-p)^{n-k}\\
P(x=1)=800p(1-p)^{800-1}=\\=800\cdot0.002\cdot0.998^{799}\approx\\
\approx0.3232=32.32\%"